bil and his brother is playing a board game . bill is currently 11 space from the finish line. if they only roll one dice per turn. what is the probability he will win in two turns
Answer by Theo(13342)

the dice has the numbers 1 through 6 on it.
in order to get to 11 or greater in 2 turns, he has to roll:

5 then 6 or 6 then 5 or 6 then 6.

the probability of any one number coming up is 1/6.

the probability of getting a 5 then a 6 is equal to 1/6 * 1/6 = 1/36.
the probability of getting a 6 then a 5 is equal to 1/6 * 1/6 = 1/36.
the probability of getting a 6 then a 6 is equal to 1/6 * 1/6 = 1/36.

the probability of getting a 5 then a 6, or a 6 then a 5, or a 6 then a 6, is equal to 1/36 + 1/36 + 1/36 = 3/36 = 1/12.